# A Host–Kra $\mathbb{F}_{2}^{\omega}$-system of order $5$ that is not Abramov of order $5$, and non-measurability of the inverse theorem for the $U^{6}(\mathbb{F}_{2}^{n})$ norm

### v2

@@ 1. Introduction
$\mathbb{E}_{x\in A}\coloneqq\frac{1}{|A|}\sum_{x\in A}$ denotes the averaging operation, and $|A|$ denotes the cardinality of a finite set $A$. If $P\colon G\to\mathbb{T}$ is a function taking values in the unit circle $\mathbb{T}\coloneqq\mathbb{R}/\mathbb{Z}$, then we have $\|e(P)\|_{U^{k+1}(G)}\leq 1$, with equality precisely when $P$ is a (non-classical) polynomial of degree $k$, as defined in Definition A.18; here $e\colon\mathbb{T}\to\mathbb{C}$ is the fundamental character $e(\theta)\coloneqq e^{2\pi i\theta}$. The space of polynomials $P\colon G\to\mathbb{T}$ of degree at most $k$ is an abelian group which we denote $\mathrm{Poly}^{k}(G)$. By convention, $\mathrm{Poly}^{0}(G)$ will denote the constant functions $\mathbb{T}$, and $\mathrm{Poly}^{k}(G)=\{0\}$ for all $k<0$ (thus non-zero constants have degree 00, and zero has degree $\infty$- ∞).
@@
In [26], a variant of the Furstenberg correspondence principle was used to show that Conjecture 1.2 implied Conjecture 1.1 for any given choice of $p,k$. In [1], Conjecture 1.2 was established in the high characteristic case $k+1\leq p$; combining the two results, this also gave Conjecture 1.1 in this regime. The full case of Conjecture 1.1 was subsequently established in [26] by a different method; alternate proofs of some or all of the cases of this conjecture have since been given in [8], [9], [4], [27], [21]. In particular Conjecture 1.2 was established in [4, Theorem 1.12] in the slightly larger range $k\leq p+1$ (and an alternate proof of Conjecture 1.1 was given for all $k,p$). We also remark that in [1, Theorem 1.20], a weaker version of Conjecture 1.2 was established in which $\mathrm{Poly}^{k}(X)$ was replaced by some unspecified subalgebra of $\mathrm{Poly}^{C(p,k)}(X)$ for some constant $C(p,k)$ depending only on $p,k$. We also note that several other structural results on ergodic $\mathbb{F}_{p}^{\omega}$-systems are known; see in particular [4], [18].
@@
Let $\eta>0$, and let $\varepsilon\colon\mathbb{N}\to\mathbb{R}^{+}$ be a decreasing function. Then there exists $M=M(k,p,\eta,\varepsilon())$ such that whenever $G=\mathbb{F}_{p}^{n}$ is an elementary abelian $p$-group and functions $f\colon G\to{\mathbb{D}}$ with $\|f\|_{U^{k+1}(G)}\geq\eta$, such that if $\vec{h}=(h_{1},\dots,h_{M})\in G^{M}$ is a tuple of shifts drawn uniformly from $G^{M}$, then with probability at least $1/2$, there exist $1\leq m\leq M$, $P\in\mathrm{Poly}^{k}(G)$ and a function $F\colon{\mathbb{D}}^{\mathbb{F}_{p}^{M}}\to\mathbb{C}$ of Lipschitz constant at most $M$ (using say the Euclidean metric on ${\mathbb{D}}^{\mathbb{F}_{p}^{M}}$), such that
@@
The numerical value of the probability $1/2$ here is inessential and could be replaced by any other constant between 00 and $1$. Roughly speaking, Conjecture 1.3 is a strengthening of Conjecture 1.1 in which the polynomial $P$ produced by that conjecture is well approximated by some combination of random shifts of $f$, where the degree $\varepsilon(m)$ of approximation can be guaranteed to be much better than the level $\frac{1}{m}$ of correlation between the polynomial $P$ and the original function $f$. The Lipschitz property of $F$ is unimportant, since one can easily discretize $f$ to take on a bounded number of values, but we retain it for minor technical reasons.
@@
When $k=1$, Conjecture 1.3 can be established by standard Fourier-analytic arguments which we now sketch here (suppressing the precise quantitative bounds needed to make precise terms such as “large” in order to simplify the exposition). If $f\colon G\to{\mathbb{D}}$ has large $U^{2}(G)$ norm, then $f$ has a large inner product with the convolution $f*f*\tilde{f}$, where $\tilde{f}(x)\coloneqq\overline{f}(-x)$. Furthermore, this convolution can be approximated in the uniform norm by a bounded linear combination of characters $e(\xi\cdot x)$. If one chooses a large number of random shifts $h_{1},\dots,h_{M}$, then with high probability one can find a convolution filter on these shifts that isolates one of these characters, that is to say there exists a linear combination $\lambda$ of the Kronecker delta functions $\delta_{h_{1}},\dots,\delta_{h_{M}}$ such that $f*f*\tilde{f}*\lambda(x)$ is close to $e(\xi\cdot x)$. By further random sampling of the $f*\tilde{f}$ factor (and increasing the number $M$ of shifts if necessary), one can with high probability approximate $f*f*\tilde{f}*\lambda$ (in $L^{1}$ norm) by a linear combination of shifts of $f$ along linear combinations of $h_{1},\dots,h_{M}$. This can then be used to establish the $k=1$ case of Conjecture 1.3; we leave the details to the interested reader.
For $k=2,3$ (and $p=2$), the strong inverse conjecture is reminiscent [^1] We are indebted to James Leng for this observation. of the quadratic Goldreich–Levin theorem from [28] (and the more recent cubic Goldreich–Levin theorem from [19]), which gives a polynomial (in $n$) time randomized algorithm to reconstruct the polynomial $P$ from the function $f$; however strong inverse conjecture is (in principle) stronger than these Goldreich–Levin type results, in that it should (after some additional effort) yield a bounded-time (rather than polynomial-time) randomized algorithm to obtain an approximation to the polynomial $P$. Such algorithms are similar in spirit [^2] We are indebted to Avi Wigderson for this remark. to implicit (or “local”) list decoding algorithms for Reed–Muller codes, as discussed for instance in [23], [7].
@@
However, the low characteristic case presents additional difficulties; for instance, a key “exact roots” property for polynomials in order $k$ $\mathbb{F}_{p}^{\omega}$-systems is known to fail in low characteristic [26, Appendix E]. In fact we are able to construct the following counterexample, which is the main result of our paper.
@@
Combining Theorem 1.6 with the contrapositive of Theorem 1.5, we conclude that Conjecture 1.2 also fails when $p=2$ and $k=5$; see also Remark 5.11 for how one might give a more direct construction of a counterexample to that conjecture. Our construction was located numerically, but we give a human-verifiable proof of the theorem here, taking advantage in particular of several technical simplifications available in the $p=2$ case (in particular, we take advantage of the ability to identify the $n$-dimensional cube $\{0,1\}^{n}$ with the $n$-dimensional vector space $\mathbb{F}_{2}^{n}$, for instance in (62)). It would be interesting to determine the complete range of $p,k$ for which Conjecture 1.3 and Conjecture 1.2 holds; for instance, the case $p=2,k=4$ remains unresolved for both conjectures, and we have not been able to rigorously establish that these conjectures are monotone [^3] There is precedent for failure of monotonicity in this subject: in [20], [21, Corollary 10] it was shown that a conjecture of Tidor [27] on approximately symmetric multilinear forms fails for $4$-linear forms but is true again for $5$-linear forms. We thank Luka Milićević for this example. in $k$.
Informally, Theorem 1.6 asserts that in characteristic two, there exist “pseudo-quintic” functions $f$ which have large $U^{6}(\mathbb{F}_{2}^{n})$ norm, and in fact correlate with a genuine quintic $e(P)$, but that the quintics that $f$ correlates with will be “non-measurable” in the sense that they cannot be approximated by a combination of boundedly many translates of $f$. Instead, one has to use “non-measurable” operations, such as taking exact roots of polynomials as in [26], in order to locate such quintics $e(P)$.
@@
Recently, quantitative versions of Conjecture 1.1 for $p=2$ and $k=3,4,5$ have been established in [27, 21]; in particular the paper [21] covers the case $p=2,k=5$ of Theorem 1.6. This is however not a contradiction; a crucial step [27, Proposition 3.5] in both those papers (a special case of Theorem 2.3 below) is the ability to represent a “strongly symmetric $k$-linear form” as the $k$-fold derivative of a degree $k$ polynomial, and this step is “non-measurable” as it requires one to expand the form into monomials using a choice of basis for $\mathbb{F}_{2}^{n}$.
@@ 1.1. Overview of proof
The function $\rho$ has to obey a certain number of properties in order to be able to find a solution $S$ to the equation (1). Firstly, $\rho$ must be symmetric with respect to permutations of $\{1,\dots,6\}$ and must also obey a certain “cocycle equation” arising from the identity $\partial_{h+k}S=\partial_{h}S+T^{h}\partial_{k}S$, where $T^{h}S(x)\coloneqq S(x+h)$ is the shift map. These properties can be formalized in the language of nilspaces by requiring $\rho$ to be a $5$-cocycle on $X_{2}$ taking values in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$; see Definition A.6 for details. However, the property of being a $5$-cocycle is not yet sufficient to guarantee a solution to (1); in the language of nilspaces, not all $5$-cocycles on $\mathbb{F}_{2}^{n}$ are $5$-coboundaries. In order to locate a solution, we will require the cocycle $\rho$ to obey an additional property that we call “strong $2$-homogeneity”. This property asserts that $\rho$ takes the form
@@
and thus we have $e(S)=e(P)e(F(Q))$ for some quintic polynomial $P\in\mathrm{Poly}^{5}(\mathbb{F}_{2}^{n})$. Morally speaking, this relation indicates that $e(P)$ correlates with $e(S)$, and that $P$ should be well approximated by $S$ and its translates (since from (1) we expect $Q$ to similarly be well approximable in this fashion).
@@
but the cube structure on $X_{5,5}$ is somewhat non-trivial, involving the cocycle $\rho\colon C^{6}(X_{2})\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ mentioned previously. Roughly speaking, the nilspace $X_{5,5}$ is the abstraction of a pair $(Q,S)$ of functions, in which $Q$ is itself a pair $Q=(Q_{1},Q_{2})$ of classical quadratic polynomials (taking values in $\mathbb{F}_{2}$), and $S$ is a “pseudo-quintic” taking values in $\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z}$ that obeys the identity (1). It will turn out not to be possible to correlate $S$ with any genuine quintics that only arise from $Q,S$, and a bounded (and randomly selected) number of their translates. The actual verification that these translates do not actually provide any useful information for the purpose of constructing a quintic turns out to be rather tricky, requiring one to show that a certain nilspace extension “splits”: see Lemma 5.7. A simpler version $X_{5,1}$ of the nilspace $X_{5,5}$, in which the cyclic group $\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z}$ is replaced by $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$, can also be used to quickly answer a question of Candela, González-Sánchez, and Szegedy [4, Question 5.18] in the negative, thus giving a weaker version of Theorem 1.6; see Proposition 4.5.
@@
We identify the core of the proof as solving a cohomological problem: finding finite abelian $2$-groups equipped with a cube structure that support $2$-homogeneous, $k$-cocycles which are not $k$-coboundaries. The cubes constructed from nilspaces built from such cocycles encode functions with large Gowers norms that violate the strong inverse Gowers conjecture 1.3.

### v3

@@ 1. Introduction
$\mathbb{E}_{x\in A}\coloneqq\frac{1}{|A|}\sum_{x\in A}$ denotes the averaging operation, and $|A|$ denotes the cardinality of a finite set $A$. If $P\colon G\to\mathbb{T}$ is a function taking values in the unit circle $\mathbb{T}\coloneqq\mathbb{R}/\mathbb{Z}$, then we have $\|e(P)\|_{U^{k+1}(G)}\leq 1$, with equality precisely when $P$ is a (non-classical) polynomial of degree $k$, as defined in Definition A.18; here $e\colon\mathbb{T}\to\mathbb{C}$ is the fundamental character $e(\theta)\coloneqq e^{2\pi i\theta}$. The space of polynomials $P\colon G\to\mathbb{T}$ of degree at most $k$ is an abelian group which we denote $\mathrm{Poly}^{k}(G)$. By convention, $\mathrm{Poly}^{0}(G)$ will denote the constant functions $\mathbb{T}$, and $\mathrm{Poly}^{k}(G)=\{0\}$ for all $k<0$ (thus non-zero constants have degree 0, and zero has degree -$\infty$).
@@
In [26], a variant of the Furstenberg correspondence principle was used to show that Conjecture 1.2 implied Conjecture 1.1 for any given choice of $p,k$. In [1], Conjecture 1.2 was established in the high characteristic case $k+1\leq p$; combining the two results, this also gave Conjecture 1.1 in this regime. The full case of Conjecture 1.1 was subsequently established in [27] by a different method; alternate proofs of some or all of the cases of this conjecture have since been given in [9], [10], [5], [28], [22]. In particular Conjecture 1.2 was established in [5, Theorem 1.12] in the slightly larger range $k\leq p+1$ (and an alternate proof of Conjecture 1.1 was given for all $k,p$). We also remark that in [1, Theorem 1.20], a weaker version of Conjecture 1.2 was established in which $\mathrm{Poly}^{k}(X)$ was replaced by some unspecified subalgebra of $\mathrm{Poly}^{C(p,k)}(X)$ for some constant $C(p,k)$ depending only on $p,k$. We also note that several other structural results on ergodic $\mathbb{F}_{p}^{\omega}$-systems are known; see in particular [5], [19].
@@
Let $\eta>0$, and let $\varepsilon\colon\mathbb{N}\to\mathbb{R}^{+}$ be a decreasing function. Then there exists $M=M(k,p,\eta,\varepsilon())$ such that the following holds. If $G=\mathbb{F}_{p}^{n}$ is an elementary abelian $p$-group and $f\colon G\to{\mathbb{D}}$ satisfies $\|f\|_{U^{k+1}(G)}\geq\eta$, then, drawing a random tuple $\vec{h}=(h_{1},\dots,h_{M})$ uniformly from $G^{M}$, we have with probability at least $1/2$ that there exist $1\leq m\leq M$, $P\in\mathrm{Poly}^{k}(G)$ and a function $F\colon{\mathbb{D}}^{\mathbb{F}_{p}^{M}}\to\mathbb{C}$ of Lipschitz constant at most $M$ (using say the Euclidean metric on ${\mathbb{D}}^{\mathbb{F}_{p}^{M}}$), such that
@@
The numerical value of the probability $1/2$ here is inessential and could be replaced by any other constant between 0 and $1$. Roughly speaking, Conjecture 1.3 is a strengthening of Conjecture 1.1 in which the polynomial $P$ produced by that conjecture is well approximated by some combination of random shifts of $f$, where the degree $\varepsilon(m)$ of approximation can be guaranteed to be much better than the level $\frac{1}{m}$ of correlation between the polynomial $P$ and the original function $f$. The Lipschitz property of $F$ is unimportant, since one can easily discretize $f$ to take on a bounded number of values, but we retain it for minor technical reasons [^1] The Lipschitz property is used in Appendix B to prove the correspondence principle (see Theorem 1.5 below), where it is used in the proof of Proposition B.1 (cf. [26, Proposition 3.13]) in the form of the Arzelà–Ascoli theorem..
@@
When $k=1$, Conjecture 1.3 can be established by standard Fourier-analytic arguments, which we now briefly sketch (suppressing precise quantitative bounds in order to simplify the exposition). If $f\colon G\to{\mathbb{D}}$ has large $U^{2}(G)$ norm, then $f$ has a large inner product with the convolution $f*f*\tilde{f}$, where $\tilde{f}(x)\coloneqq\overline{f}(-x)$. Furthermore, this convolution can be uniformly approximated by a bounded linear combination of characters $e(\xi\cdot x)$, which is standard. To isolate one such frequency $\xi$, one may choose a large number of random shifts $h_{1},\dots,h_{M}$ and then, with high probability, construct a single linear combination $\lambda$ of the delta masses $\delta_{h_{1}},\dots,\delta_{h_{M}}$ such that $f*f*\tilde{f}*\lambda$ is uniformly [^2] Take $\xi_{0}$ to be a frequency with $|\widehat{f}(\xi_{0})|$ maximal, and define $\lambda=\frac{1}{Mc_{0}}\sum_{j=1}^{M}\overline{e(\xi_{0}\cdot h_{j})}\delta_{-h_{j}},\qquad c_{0}=\widehat{(f*f*\tilde{f})}(\xi_{0}),$ for i.i.d. uniform shifts $h_{j}$. Then $\widehat{\lambda}(\xi_{0})=1/c_{0}$ exactly, while for every other significant frequency $\xi\neq\xi_{0}$, the value $\widehat{\lambda}(\xi)=\tfrac{1}{Mc_{0}}\sum_{j}e((\xi-\xi_{0})\cdot h_{j})$ is an average of mean-zero, unit-modulus i.i.d. variables. By Hoeffding’s inequality and a union bound over the finitely many “large” Fourier coefficients, these averages are simultaneously small with high probability, so $f*f*\tilde{f}*\lambda$ is uniformly close to the pure character $x\mapsto e(\xi_{0}\cdot x)$. close to the character $x\mapsto e(\xi\cdot x)$. Finally, by further random sampling of the $f*\tilde{f}$ factor (and increasing $M$ if necessary), one can approximate $f*f*\tilde{f}*\lambda$ in $L^{1}$ by a bounded linear combination of shifts of $f$ along linear combinations of $h_{1},\dots,h_{M}$. This yields the case $k=1$ of Conjecture 1.3; we leave the details to the interested reader.
For $k=2,3$ (and $p=2$), the strong inverse conjecture is reminiscent [^3] We are indebted to James Leng for this observation. of the quadratic Goldreich–Levin theorem from [29] (and the more recent cubic Goldreich–Levin theorem from [20]), which gives a polynomial (in $n$) time randomized algorithm to reconstruct the polynomial $P$ from the function $f$; however, the strong inverse conjecture is (in principle) stronger than these Goldreich–Levin type results, in that it should (after some additional effort) yield a bounded-time (rather than polynomial-time) randomized algorithm to obtain an approximation to the polynomial $P$. Such algorithms are similar in spirit [^4] We are indebted to Avi Wigderson for this remark. to implicit (or “local”) list decoding algorithms for Reed–Muller codes, as discussed for instance in [24], [8].
@@
However, the low characteristic case presents additional difficulties; for instance, a key “exact roots” property for polynomials in order $k$ $\mathbb{F}_{p}^{\omega}$-systems is known to fail in low characteristic [27, Appendix E]. In fact we are able to construct the following counterexample, which is the main result of our paper.
@@
Combining Theorem 1.6 with the contrapositive of Theorem 1.5, we conclude that Conjecture 1.2 also fails when $p=2$ and $k=5$; see also Remark 5.11 for how one might give a more direct construction of a counterexample to that conjecture. Our construction was located numerically, but we give a human-verifiable proof of the theorem here, taking advantage in particular of several technical simplifications available in the $p=2$ case (in particular, we take advantage of the ability to identify the $n$-dimensional cube $\{0,1\}^{n}$ with the $n$-dimensional vector space $\mathbb{F}_{2}^{n}$, for instance in (65)). It would be interesting to determine the complete range of $p,k$ for which Conjecture 1.3 and Conjecture 1.2 holds; for instance, the case $p=2,k=4$ remains unresolved for both conjectures, and we have not been able to rigorously establish that these conjectures are monotone [^5] There is precedent for failure of monotonicity in this subject: in [21], [22, Corollary 10] it was shown that a conjecture of Tidor [28] on approximately symmetric multilinear forms fails for $4$-linear forms but is true again for $5$-linear forms. We thank Luka Milićević for this example. in $k$.
Informally, Theorem 1.6 asserts that in characteristic two, there exist “pseudo-quintic” functions $f$ which have large $U^{6}(\mathbb{F}_{2}^{n})$ norm, and in fact correlate with a genuine quintic $e(P)$, but that the quintics that $f$ correlates with will be “non-measurable” in the sense that they cannot be approximated [^6] To justify this intuition, observe from the Stone–Weierstrass theorem that if $\mathcal{F}$} is the $\sigma$-algebra generated by a function $f\in L^{\infty}(X)$ and its $G$-translates, then a function $g\in L^{\infty}(X)$ will be $\mathcal{F}$}-measurable iff it can approximated in $L^{1}$ to arbitrary accuracy by a polynomial combination of a finite number of translates of $f$ and $\overline{f}$. To “finitize” this assertion by replacing the measure-preserving system $X$ with a finite space $\mathbb{F}_{2}^{n}$, we informally replace “finite number of translates” with “bounded number of translates”, to prevent the notion of “measurability” from becoming trivial in the finitary setting. in $L^{1}$ by a polynomial combination of boundedly many translates of $f$ and its complex conjugate. Instead, one has to use “non-measurable” operations, such as taking exact roots of polynomials as in [27], in order to locate such quintics $e(P)$.
@@
Recently, quantitative versions of Conjecture 1.1 for $p=2$ and $k=3,4,5$ have been established in [28, 22]; in particular the paper [22] covers the case $p=2,k=5$ of Theorem 1.6. This is however not a contradiction; a crucial step [28, Proposition 3.5] in both those papers (a special case of Theorem 2.3 below) is the ability to represent a “strongly symmetric $k$-linear form” as the $k$-fold derivative of a degree $k$ polynomial, and this step is “non-measurable” as it requires one to expand the form into monomials using a choice of basis for $\mathbb{F}_{2}^{n}$.
@@ 1.1. Overview of proof
The function $\rho$ has to obey a certain number of properties in order to be able to find a solution $S$ to the equation (1). Firstly, $\rho$ must be symmetric with respect to permutations of $\{1,\dots,6\}$ and must also obey a certain “cocycle equation” arising from the identity $\partial_{h+k}S=\partial_{h}S+T^{h}\partial_{k}S$, where $T^{h}S(x)\coloneqq S(x+h)$ is the shift map. These properties can be formalized in the language of nilspaces by requiring $\rho$ to be a $5$-cocycle on $X_{2}$ taking values in $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$; see Definition A.6 for details. However, the property of being a $5$-cocycle is not yet sufficient to guarantee a solution to (1). In order to locate a solution, we will require the cocycle $\rho$ to obey an additional property that we call “strong $2$-homogeneity”. This property asserts that $\rho$ takes the form
@@
and thus we have $e(S)=e(P)e(F(Q))$ for some quintic polynomial $P\in\mathrm{Poly}^{5}(\mathbb{F}_{2}^{n})$. Morally speaking, this relation indicates that $e(P)$ correlates with $e(S)$, and that $P$ should be well approximated by $S$ and its translates. Indeed, note that from the identity e(S)=e(P)e(F(Q))e(S)=e(P)\,e(F(Q)) it follows that the correlation of $e(S)$ with $e(P)$ is simply the expectation of $e(F(Q(x)))$, which we expect to be nontrivial since $Q$ is a quadratic. Moreover, from the relation $e(P)=e(S)\,\overline{e(F(Q(x)))}$ together with (1), one sees that one can approximate $P$ by combining information from $S$ and its shifts.
@@
but the cube structure on $X_{5,5}$ is somewhat non-trivial, involving the cocycle $\rho\colon C^{6}(X_{2})\to\frac{1}{2}\mathbb{Z}/\mathbb{Z}$ mentioned previously. Roughly speaking, the nilspace $X_{5,5}$ is the abstraction of a pair $(Q,S)$ of functions, in which $Q$ is itself a pair $Q=(Q_{1},Q_{2})$ of classical quadratic polynomials (taking values in $\mathbb{F}_{2}$), and $S$ is a “pseudo-quintic” taking values in $\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z}$ that obeys the identity (1). It will turn out not to be possible to correlate $S$ with any genuine quintics that only arise from $Q,S$, and a bounded (and randomly selected) number of their translates. The actual verification that these translates do not actually provide any useful information for the purpose of constructing a quintic turns out to be rather tricky, requiring one to show that a certain nilspace extension “splits”: see Lemma 5.7. A simpler version $X_{5,1}$ of the nilspace $X_{5,5}$, in which the cyclic group $\frac{1}{2^{5}}\mathbb{Z}/\mathbb{Z}$ is replaced by $\frac{1}{2}\mathbb{Z}/\mathbb{Z}$, can also be used to quickly answer a question of Candela, González-Sánchez, and Szegedy [5, Question 5.18] in the negative, thus giving a weaker version of Theorem 1.6; see Proposition 4.5.
@@
We identify the core of the proof as solving a cohomological problem: finding finite abelian $2$-groups equipped with a cube structure that support $2$-homogeneous $k$-cocycles which are not $k$-coboundaries. The cubes constructed from nilspaces built from such cocycles encode functions with large Gowers norms that violate the strong inverse Gowers conjecture 1.3.